# Quantum cryptography beyond key distribution: theory and experiment

### v1

@@ I.1 General motivation
Crucially, the rise of quantum algorithms providing super-polynomial speed-up for solving these tasks has triggered a response to search for stronger security levels Shor (1997); Martín-López et al. (2012); Gouzien and Sangouard (2021). Two main approaches are being actively developed: the first approach uses classical resources to design primitives that are secure against quantum attacks. Labeled post-quantum cryptography, this field is evolving very rapidly, both in terms of standardization and cryptanalysis, in which security conjectures are broken and new schemes appear. The second approach, which is the main focus of this review, employs quantum resources to protect against quantum attacks, and is typically referred to as quantum cryptography.
@@ I.2 Aims, scope and structure
We follow the vision of the initial review on Quantum cryptography beyond quantum key distribution Broadbent and Schaffner (2016) and the preliminary steps taken by the Quantum Protocol Zoo qzo (2024) in covering literature outside the widely studied field of QKD and quantum random number generation. For QKD, we refer the Reader to previous reviews discussing its theoretical aspects Portmann and Renner (2022) and practical challenges Xu et al. (2020a); Gisin et al. (2002); Pirandola et al. (2020); Diamanti et al. (2016). For quantum random number generation, we refer the Reader to Ma et al. (2016); Herrero-Collantes and Garcia-Escartin (2017).
In Section II, we recall the fundamental principles of quantum cryptography and some mathematical tools that will be useful throughout the review. In particular, we provide some intuition about the search for optimal cheating strategies, which is a core component of many quantum-cryptographic security proofs.
@@ II.4 Optimal cheating strategies
Proving the security of quantum primitives in the i.t.-secure setting often translates to the search for the optimal cheating strategy. If no explicit strategy can be found, one should at least derive an upper bound on its success probability. To achieve this goal, the high-level guidelines in this section are based on the following assumptions:
@@ II.6 Minimal assumptions
In the classical world, many cryptographic primitives can be achieved securely under the assumption that one-way functions (OWFs), namely functions that can be evaluated easily but are computationally hard to invert, exist Impagliazzo and Luby (1989). In the quantum world however, it turns out that many such primitives can be securely implemented under even weaker assumptions.
@@ III.1 Unforgeable tokens
The most intuitive cryptographic application of quantum no-cloning was arguably formulated in 1973, and published ten years later Wiesner (1983). In this pioneering proposal, conjugate coding was introduced as a means to generate quantum objects that are inherently impossible to counterfeit, such as quantum banknotes. Note that, in the classical world, there exists no such notion of security for banknotes. Instead, the difficulty of producing counterfeits relies on their intricate coloring and watermark design Chia and Levene (2009).
@@
Perhaps the most significant challenge for practical quantum tokens resides in the need for efficient, high fidelity quantum storage as a means to ensure flexible spending. Quantum storage times of a few hours or days are currently beyond state-of-the-art, which ranges from a few microseconds to a few minutes Ma et al. (2021); Vernaz-Gris et al. (2018); Heshami et al. (2016). Recently however, interesting alternatives were proposed, replacing quantum storage by a network of trusted agents and authenticated channels, positioned at precise spatial locations with respect to the verification points Kent and Pitalúa-García (2020); Kent (2019). To some extent, these schemes allow the user to choose the time and place of verification without storing the quantum states. From a security perspective, monitoring the spatial and temporal coordinates of verifiers requires either a trusted Global Positioning System (GPS) or secure position verification schemes discussed in Section III.3. A first partial implementation using weak coherent states was performed in Kent et al. (2022), followed by an experiment demonstrating transaction time advantage over classical schemes Jiang et al. (2024). Note that the work from Kent et al. (2022) also stated three properties that should be simulateneously achieved to claim a quantum advantage: token unforgeability, instant validation and user privacy.
@@ III.2 Unclonable encryption
Although this first quantum proposal allowed for tamper detection, the question of how much information can be extracted from the original message by splitting the ciphertext into several (imperfect) quantum copies was left open. This was addressed by introducing the somewhat stronger notion of unclonable quantum encryption, whose security guarantees may be understood as a three-partite game between an honest sender Alice and two dishonest receivers Bob and Charlie Broadbent and Lord (2020). Essentially, Alice chooses a message $m$ of length $l$ and an encryption key $k$ of length polynomial in $\lambda$ (a security parameter), producing a quantum ciphertext $\sigma\left(m,k\right)$ on $\mathcal{H}_{A}$. Following the intuition from Section II.4, Bob and Charlie apply some optimal CPTD map $\Lambda:\mathcal{H}_{A}\rightarrow\mathcal{H}_{B}\otimes\mathcal{H}_{C}$, each keeping register $B$ and $C$, respectively, before being isolated from one another. In a later phase, Alice reveals $k$ to both of them, allowing them to recover two messages $m_{B}$ and $m_{C}$ from their respective quantum systems. Bob and Charlie win the game if and only if their messagea agree with the originale one (i.e. $m_{B}=m_{C}=m$). Following Broadbent and Lord (2020), the scheme is $t$-unclonable if their winning probability is upper bounded by $2^{-l+t}+negl(\lambda)$, where $negl(\lambda)$ indicates a negligible parameter in $\lambda$.
@@ III.4 Covert communication
In 2015, i.t-secure bounds were established for covert communication over lossy thermal-noise bosonic channels Bash et al. (2015), which quantum-mechanically describe the propagation of a single spatio-temporal mode of the electromagnetic field through a lossy channel with thermal noise. Assuming that Alice and Bob pre-share a long enough secret, and that there exists some excess noise that cannot be controlled by the quantum adversary –eg: blackbody radiation–, then Alice can covertly transmit $O\left(\sqrt{n}\right)$ bits to Bob using $n$ photonic modes Bullock et al. (2020).
@@
Possibilities of performing covert QKD for the re-generation of pre-shared secrets was investigated, first deemed impossible with i.t.-security Arrazola and Scarani (2016), but later revisited using more sophisticated coding schemes for information reconciliation and privacy amplification Tahmasbi and Bloch (2019). For completeness, we note that security bounds on steganography with quantum information were also derived Shaw and Brun (2011); Sutherland and Brun (2019).
@@ III.5 Fingerprinting
One important communication task involves two parties comparing their respective bit strings with as little communication and resources as possible, in order to determine whether they are equal. More formally, Alice and Bob each hold an $n$-bit string, labelled $x$ and $y$ respectively, and they wish to check whether $x=y$ or not. The protocol occurs in the setting of simultaneous message passing [^2] The simultaneous message passing model has many applications in complexity theory, cryptography, distributed systems, IoT, machine learning, and multi-agent systems, where it is essential to reduce direct communication rounds among parties while still achieving computational goals., where (a) each party sends their message to a common referee who will determine whether they are identical and (b) the parties do not share any common randomness source (or entanglement). Classically, this task requires each player to send $\Omega(\sqrt{n})$-bit messages Newman and Szegedy (1996).
@@ IV.1 Digital signatures
In parallel to these theoretical developments, we discuss the key experimental progress related to QDS. The first proof-of-principle experiment used phase-encoded coherent states Clarke et al. (2012), allowing each participant to split their state in two and compare each half with another participants’ half through interference on a shared beamsplitter, along with an unambiguous state discrimination (USD) measurement. The setting-up and signing phases were performed one after the other to avoid the use of a quantum memory. A protocol that did not require any quantum storage was later implemented Collins et al. (2014), removing the need to sign immediately after the setting-up phase. An unambiguous state elimination measurement (generalization of the USD measurement used earlier) was implemented. The first demonstration of QDS in the continuous variable setting was performed in Croal et al. (2016), using free-space channels and heterodyne detection for the measurements. Experiments exceeding the kilometer range were then achieved Donaldson et al. (2016), followed by measurement-device independent schemes Yin et al. (2017); Roberts et al. (2017). An i.t.-secure implementation over an eight-user quantum network was also provided Pelet et al. (2022).
@@ IV.2 Bit commitment
Under such computational assumptions, BC comes in two dual flavors: binding but computationally hiding and hiding but computationally binding. Commitments of the first type may be achieved from any one-way function Naor (1991); Håstad et al. (1999). The construction of computationally binding and statistically hiding BC from any one-way function was solved more recently in Haitner et al. (2009).
@@
Remarkably, values of $P_{s}$ and $P_{r}$ arbitrarily close to $1/2$ may be achieved using additional physical assumptions such as no-signalling constraints or the fact that an adversary’s quantum storage is imperfect or bounded. These schemes, along with their experimental demonstrations, are covered in Section V.
@@ IV.3 Oblivious transfer
The quantum technology necessary to implement QOT is essentially the same as QKD. The initial prototype of Bennett et al. (1992) had its software adapted by Bennett et al. (1991) to implement QOT. The main difference from QKD was the use of non-interactive error-correction. Interactive error-correction, performed in the original QKD, was not appropriate to the distrustful two-party scenario. Modern composable security definitions have encouraged the development of more advanced QOT implementations as proposed in Diamanti et al. (2024).

### v2

@@ I.1 General motivation
Crucially, the rise of quantum algorithms providing super-polynomial speed-up for solving these tasks has triggered a response to search for stronger security levels Shor (1997); Gouzien and Sangouard (2021); Gidney (2025). Two main approaches are being actively developed: the first approach uses classical resources to design primitives that are secure against quantum attacks. Labeled post-quantum cryptography, this field is evolving very rapidly, both in terms of standardization and cryptanalysis, in which security conjectures are broken and new schemes appear. The second approach, which is the main focus of this review, employs quantum resources to protect against quantum attacks, and is typically referred to as quantum cryptography.
@@ I.2 Aims, scope and structure
We follow the vision of the initial review on Quantum cryptography beyond quantum key distribution Broadbent and Schaffner (2016) and the preliminary steps taken by the Quantum Protocol Zoo qzo (2024) in covering literature outside the widely studied field of QKD and quantum random number generation. In addition to Broadbent and Schaffner (2016), we survey the rapidly expanding theoretical progress of the $9$ last years, together with practical security proofs, experimental demonstrations and hardware challenges for the first time. These are particularly crucial to the development of secure quantum networks and large-scale deployment strategies.
For the Reader interested in QKD, we reference some existing reviews discussing its theoretical aspects Portmann and Renner (2022) and practical challenges Xu et al. (2020a); Gisin et al. (2002); Pirandola et al. (2020); Diamanti et al. (2016). For quantum random number generation, we refer the Reader to Ma et al. (2016); Herrero-Collantes and Garcia-Escartin (2017).
Our review is structured as follows: In Section II, we recall the fundamental principles of quantum cryptography and some mathematical tools that will be useful throughout the review. In particular, we provide some intuition about the search for optimal cheating strategies, which is a core component of many quantum-cryptographic security proofs.
@@ II.4 Optimal cheating strategies
Proving the security of quantum primitives in the i.t.-secure setting often translates to the search for the attacker’s optimal cheating strategy. If no explicit strategy can be found, one should at least derive an upper bound on its success probability, which quantifies either how likely it is that the protocol fails to prevent the attack in the first place or how likely it is that a successful attack remains undetected by the honest part(ies). To achieve this goal, the high-level guidelines in this section are based on the following assumptions:
@@ II.6 Minimal assumptions
As summarized in Table 1, some primitives exploiting quantum resources can achieve i.t.-security under various physical assumptions, while others may only achieve computational security against quantum attacks. In the latter setting, it is important to identify which are the minimal or weakest assumptions required to achieve a given primitive, some of which are introduced in this section.
In the classical world, many primitives can be achieved securely under the assumption that one-way functions (OWFs), namely functions that can be evaluated easily but are computationally hard to invert, exist Impagliazzo and Luby (1989). In the quantum world however, it turns out that many such primitives can be securely implemented under even weaker assumptions.
@@
Security Assumption Primitive
information -theoretic quantum mechanics (QM) unforgeable tokens, unclonable encryption, data locking, digital signatures, weak coin flipping, cheat-sensitive imperfect mistrustful primitives (Section IV), multipartite primitives (Section VII), private quantum channels, authentication of quantum messages, anonymous communication, blind computation with quantum clients
QM + noisy channel covert communication
QM + special relativity bit commitment, oblivious transfer, strong coin flipping, multiparty computation
QM + bounded/noisy quantum storage data-locked QKD, bit commitment, oblivious transfer, strong coin flipping, multiparty computation
QM + special relativity + bounded entanglement position verification
computational pseudorandom quantum states, one-way state generators, EFI pairs bit commitment, oblivious transfer, strong coin flipping, zero-knowledge proofs
quantum-safe one-way functions encryption with quantum public keys
trapdoor $2$-regular collision-resistant, trapdoor claw-free functions blind computation with classical clients, fully homomorphic encryption
indistinguishability obfuscation public-key money
indistinguishability obfuscation + learning with errors one-shot signatures
hardware-based one-time memories one-time programs
quantum physical unclonable functions device authentication, identification, software licensing
Table 1: Classification of quantum primitives according to their achievable security guarantees –either information-theoretic, computational or hardware-based– and the assumptions required for these to hold. For detailed discussions, please refer to each primitive’s dedicated section.
@@ III.1 Unforgeable tokens
The most intuitive cryptographic application of quantum no-cloning was arguably formulated in 1973, and published ten years later Wiesner (1983). In this pioneering proposal, conjugate coding was introduced as a means to generate quantum objects that are inherently impossible to counterfeit, such as quantum banknotes. From one valid banknote associated with a given serial number, we wish that the probability of producing $k\geqslant 2$ valid copies with the same serial number can be made arbitrarily small in some security parameter. Note that, in the classical world, there exists no such notion of security for banknotes. Instead, the difficulty of producing counterfeits relies on their intricate coloring and watermark design Chia and Levene (2009).
@@
Perhaps the most significant challenge for practical quantum tokens resides in the need for efficient, high fidelity quantum storage as a means to ensure flexible spending. Quantum storage times of a few hours or days are currently beyond state-of-the-art, which ranges from a few microseconds to a few minutes Ma et al. (2021); Vernaz-Gris et al. (2018); Heshami et al. (2016). Nevertheless, two experimental demonstrations of secure quantum tokens with genuine quantum storage were realized, the first using a cold-atom-based quantum memory Mamann et al. (2025) and the second employing color centers in diamond Strocka et al. (2025).
Memoryless alternatives were also proposed, achieving spending flexibility through a network of trusted agents and authenticated channels, positioned at precise spatial locations with respect to the verification points Kent and Pitalúa-García (2020); Kent (2019). To some extent, these schemes allow the user to choose the time and place of verification without storing the quantum states. From a security perspective, monitoring the spatial and temporal coordinates of verifiers requires either a trusted Global Positioning System (GPS) or secure position verification schemes discussed in Section III.3. A first partial implementation using weak coherent states was performed in Kent et al. (2022), followed by an experiment demonstrating transaction time advantage over classical schemes Jiang et al. (2024). Note that the work from Kent et al. (2022) also stated three properties that should be simulateneously achieved to claim a quantum advantage: token unforgeability, instant validation and user privacy.
@@ III.2 Unclonable encryption
Although this first quantum proposal allowed for tamper detection, the question of how much information can be extracted from the original message by splitting the ciphertext into several (imperfect) quantum copies was left open. This was addressed by introducing the somewhat stronger notion of unclonable quantum encryption, whose security guarantees may be understood as a three-partite game between an honest sender Alice and two dishonest receivers Bob and Charlie Broadbent and Lord (2020). Essentially, Alice chooses a message $m$ of length $l$ and an encryption key $k$ of length polynomial in $\lambda$ (a security parameter), producing a quantum ciphertext $\sigma\left(m,k\right)$ on $\mathcal{H}_{A}$. Following the intuition from Section II.4, Bob and Charlie apply some optimal CPTD map $\Lambda:\mathcal{H}_{A}\rightarrow\mathcal{H}_{B}\otimes\mathcal{H}_{C}$, each keeping register $B$ and $C$, respectively, before being isolated from one another. In a later phase, Alice reveals $k$ to both of them, allowing them to recover two messages $m_{B}$ and $m_{C}$ from their respective quantum systems. Bob and Charlie win the game if and only if their messages agree with the original one (i.e. $m_{B}=m_{C}=m$). Following Broadbent and Lord (2020), the scheme is $t$-unclonable if their winning probability is upper bounded by $2^{-l+t}+negl(\lambda)$, where $negl(\lambda)$ indicates a negligible parameter in $\lambda$.
@@ III.4 Covert communication
Using quantum resources, covertness may be achieved and quantified in the context of measurement theory: security is guaranteed if a quantum adversary cannot reliably distinguish the quantum state they intercept when the parties communicate from the quantum state they intercept when the parties do not communicate. In 2015, i.t-secure bounds were established for covert communication over lossy thermal-noise bosonic channels Bash et al. (2015), which quantum-mechanically describe the propagation of a single spatio-temporal mode of the electromagnetic field through a lossy channel with thermal noise. Assuming that Alice and Bob pre-share a long enough secret, and that there exists some excess noise that cannot be controlled by the quantum adversary –eg: blackbody radiation–, then Alice can covertly transmit $O\left(\sqrt{n}\right)$ bits to Bob using $n$ photonic modes Bullock et al. (2020).
@@
Possibilities of performing covert QKD for the re-generation of pre-shared secrets was investigated, first deemed impossible with i.t.-security Arrazola and Scarani (2016), but later revisited using more sophisticated coding schemes for information reconciliation and privacy amplification Tahmasbi and Bloch (2019). Notably, schemes allowing for randomness (or secret key) expansion were designed, based on the observation that a secure covert protocol implies a secure message transmission protocol. In other words, starting from an input secret, a new secret can be communicated to the other party Arrazola and Amiri (2018). For completeness, we note that security bounds on steganography with quantum information were also derived Shaw and Brun (2011); Sutherland and Brun (2019).
Finally, using quantum entanglement, a special class of covert protocols enabling the undetectable sensing of a target phase was demonstrated Hao et al. (2022), inspired by the covert sensing schemes from Gagatsos et al. (2019); Bash et al. (2017). A little later, covert sensing designed for the Doppler effect Hao and Zhang (2023) and target detection Tham et al. (2024) were developed.
@@ III.5 Fingerprinting
One important communication task involves two parties comparing their respective bit strings with as little communication and resources as possible, in order to determine whether they are equal. More formally, Alice and Bob each hold an $n$-bit string, labelled $x$ and $y$ respectively, and they wish to check whether $x=y$ or not. The protocol occurs in the setting of simultaneous message passing [^2] The simultaneous message passing model has many applications in complexity theory, cryptography, distributed systems, IoT, machine learning, and multi-agent systems, where it is essential to reduce direct communication rounds among parties while still achieving computational goals., where (a) each party sends their message to a common referee who will determine whether they are identical and (b) the parties do not share any common randomness source (or entanglement).
Developing fingerprinting schemes that require as little communication as possible is relevant in many areas of cryptography, including secure identification protocols (eg: password-based identification in which a party wishes to authenticate themselves to a device without revealing their actual password) and secure two-party computation in the context of equality testing Couteau (2018) and the socialist millionaire problem Jakobsson and Yung (1996). Classically, it was shown that fingerprinting requires each player to send $\Omega(\sqrt{n})$-bit messages Newman and Szegedy (1996).
@@
III.6 Data locking
In a classical setting, consider Alice encrypting an $m$-bit message into an $m$-bit ciphertext using a $k$-bit key. If Alice sends the ciphertext to Bob without the key, the parties share a mutual information of $m$ bits. Once Alice communicates the key to Bob, their mutual information increases to $(m+k)$ bits. Generally, classical encryption can made i.t-secure if and only if the secret has the same size as the message, i.e. when $k=m$ Miller (1882); Vernam (1926).
Remarkably, quantum mechanics allows communication of a $k$-bit key to increase the mutual information between Alice and Bob by an amount strictly larger than $k$ bits DiVincenzo et al. (2004), which is in stark contrast to classical one-time pad encryption. Thus, it is possible to lock a uniformly random $m$-bit message into a quantum state with a key of length $k<<m$ Dupuis et al. (2013). Only once the key is revealed can the full $m$-bit data be unlocked. In fact, there exist quantum locking schemes in which the key size is independent of the message length: $n$ bits can be encoded into $n$ qubits using a secret of constant size Fawzi et al. (2013). On the other hand, this suggests that any small leakage of the secret key may unlock a large amount of information to an eavesdropper Lupo et al. (2014), which calls for protocols that are fault-tolerant and robust against such leakage Huang et al. (2021).
More generally, like many primitives discussed in Section X.1, the security definition of data locking should be carefully formulated to ensure its composability König et al. (2007). In the context of quantum-locked key distribution for example, the resulting secret key can be securely used for one-time pad encryption if the eavesdropper’s quantum storage is limited in time Lupo and Lloyd (2014). A more thorough discussion of i.t. security based on bounded and noisy storage assumptions is provided in Section V.2.
In the presence of noise and losses Fawzi et al. (2013), the performance of locking protocols was analyzed by formally defining the locking capacity of a quantum channel Guha et al. (2014). Photonic demonstrations were provided for the scheme of DiVincenzo et al. (2004), in which the eavesdropper’s accessible information is limited to half that of the legitimate receiver, and for the loss-tolerant protocol from Fawzi et al. (2013). The first scheme requires the combined channel transmission and detection efficiency to exceed $50\%$, similarly to some trustful primitives in Sections III.1 and III.3. This was achieved by employing heralded spontaneous parametric down-conversion emitting orthogonally-polarized photon pairs at telecommunication wavelengths, where Alice’s heralding employs a superconducting nanowire detector and Bob’s detection employs highly-efficient transition edge sensors Liu et al. (2016). The second, loss-tolerant scheme was implemented using the same setup with active polarization encoding and decoding through Pockels cells, surpassing the performance of classical one-time pad encryption for various values of channel loss Liu et al. (2016).
Generally, data locking has tight links to quantum secure direct communication (QSDC), in which messages are efficiently encoded onto quantum states without prior sharing of a secret key, enabling either the detection of an eavesdropper before information is leaked Long and Liu (2002); Deng and Long (2004); Wang et al. (2005) or denying access to the information through the use of entangled pairs Deng et al. (2003); Schumacher and Westmoreland (2006); Zheng and Long (2014). Along with the first QSDC experiment using frequency coding Hu et al. (2016), a quantum data locking implementation targeted to QSDC was provided, where $6$ bits of information were encoded onto a single photon using a key of less than $6$ bits Lum et al. (2016). While this work also generates photon pairs through parametric down conversion, Alice’s phase-scrambling unitaries are implemented through spatial light modulation, and the states are detected on Bob’s single-photon detection nanowire array.
A little later, QSDC employing an atomic quantum memory was reported Zhang et al. (2017) and long-distance experiments were provided Zhu et al. (2017); Zhang et al. (2022), along with continuous-variable implementations Cao et al. (2023). For a more complete review of QSDC and its experimental demonstrations, we refer the Reader to Pan et al. (2024).
@@ IV.1 Digital signatures
In parallel to these theoretical developments, we discuss the key experimental progress related to QDS. The first proof-of-principle experiment used phase-encoded coherent states Clarke et al. (2012), allowing each participant to split their state in two and compare each half with another participants’ half through interference on a shared beamsplitter, along with an unambiguous state discrimination (USD) measurement. The setting-up and signing phases were performed one after the other to avoid the use of a quantum memory. A protocol that did not require any quantum storage was later implemented Collins et al. (2014), removing the need to sign immediately after the setting-up phase. An unambiguous state elimination measurement (generalization of the USD measurement used earlier) was implemented. The first demonstration of QDS in the continuous variable setting was performed in Croal et al. (2016), using free-space channels and heterodyne detection for the measurements. Experiments exceeding the kilometer range were then achieved Donaldson et al. (2016), followed by measurement-device independent schemes Yin et al. (2017); Roberts et al. (2017). More recently, an i.t.-secure implementation over an eight-user quantum network was provided Pelet et al. (2022), along with efficient QDS based on one-time universal hashing Yin et al. (2023); Li et al. (2023a) and their application to e-commerce scenarios Cao et al. (2024).
@@ IV.2 Bit commitment
Under such computational assumptions, BC comes in two dual flavors: binding but computationally hiding and hiding but computationally binding. Commitments of the first type may be achieved from any one-way function Naor (1991); doi:10.1137/S0097539793244708. The construction of computationally binding and statistically hiding BC from any one-way function was solved more recently in doi:10.1137/080725404.
@@
Remarkably, values of $P_{s}$ and $P_{r}$ arbitrarily close to $1/2$ may be achieved using additional physical assumptions such as no-signalling constraints or the fact that an adversary’s quantum storage is imperfect or bounded. These schemes, along with their experimental demonstrations, are covered in Section V.1.
@@ IV.3 Oblivious transfer
The quantum technology necessary to implement QOT is essentially the same as QKD. The initial prototype of Bennett et al. (1992) had its software adapted by Bennett et al. (1991) to implement QOT. The main difference from QKD was the use of non-interactive error-correction. Interactive error-correction, performed in the original QKD, was not appropriate to the distrustful two-party scenario. Along with experimental demonstrations of QOT assuming only symmetric primitives to implement commitments Lemus et al. (2025), modern composable security definitions have encouraged the development of more advanced QOT implementations as proposed in Diamanti et al. (2024).
@@
For completeness, we note the existence of a related primitive known as quantum private queries Giovannetti et al. (2008); Jakobi et al. (2011), in which a user wishes to access an entry from a database while keeping their choice of entry private from the database provider. While secure $1$-out-of-$n$ OT could provide such a service, security can be achieved in a weaker cheat-sensitive model in which the provider does not wish to risk being detected cheating. A fault-tolerant protocol along with its proof-of-principle demonstration were provided in Chan et al. (2014).